OpenStudy (anonymous):

Simplify squar root of 8* square root of 10 a. 4*sq rt of 5 b. 16* sq rt of 5 c. 10* sq rt of 2 d. 40* sq rt of 2

OpenStudy (anonymous):

A ?

OpenStudy (unklerhaukus):

\[\sqrt 8\times\sqrt{10}\] \[=\sqrt{4\times2}{\times\sqrt{5\times2}}\]

OpenStudy (zzr0ck3r):

sqrt(10)*sqrt(8) = sqrt(80) = sqrt(20*4) = sqrt(4*5*4) = 2sqrt(4*5) = 4sqrt(5)

OpenStudy (zzr0ck3r):

yes A